OpenStudy (anonymous):

solve the equation below for x. 4^x=5.6962

OpenStudy (anonymous):

round to 2 decimal places

OpenStudy (anonymous):

I got 1.25? or 1.26 I think

OpenStudy (unklerhaukus):

take the logarithm of base 4 to both side of the equation

OpenStudy (anonymous):

yeah that's how I got my answer

OpenStudy (anonymous):

1.25 ?

OpenStudy (unklerhaukus):

Yep

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

yes, I just wanted to make sure because my teacher wants us to be very specific on rounding, where I got 1.25 and then 1.26

OpenStudy (anonymous):

so I wasn't sure :p

OpenStudy (unklerhaukus):

\[4^x=5.6962\\ x=\log_4(5.6962)=\frac{\ln(5.6962)}{\ln(4)}=1.254999900921=1.25 (\text{2dps})\]

OpenStudy (anonymous):

thank u both :)